# Open Packing in Graphs: Bounds and Complexity

### v1

@@ 4.2 $I_{r}$-split graphs
Figure 9: (a) An instance $(X_{1},X_{2},X_{3},X_{4},M)$ of a $4$-Dimensional Matching and (b) An $I_{r}$-split graph $G(C\cup I,E)$ corresponding to the instance $(X_{1},X_{2},X_{3},X_{4},M)$ using Construction 4. The edges between the vertices of $C$ aren’t given in the figure for clarity. Note that for every $X_{i}=\{z_{i},u_{i},v_{i}\}$ in the instance of $4$-Dimensional Matching, three vertices $z_{(x_{i},i)},z_{(u_{i},i)}$ and $z_{(v_{i},i)}$ are created in the graph $G$.
An example of Construction 4 is given in Fig. 9. The vertex and edge sets of the graph $G$ are $V(G)=C\cup I$, where $I=\{y_{w}\,:\,w\in M\}$ and $E(G)=\{xx^{\prime}\,:\,x,x^{\prime}\in C\}\cup\{z_{(x,i)}y_{w}\,:\,x\in X_{i},w\in M\text{ and the }i^{th}\text{ coordinate of }w\text{ is }x\}$. Then, $|V(G)|=(r\cdot q)+|M|$ and $|E(G)|=r\cdot|M|+\binom{r\cdot q}{2}$. Hence, the graph $G$ can be constructed in polynomial time with respect to $r,q$ and $|M|$. Note that by the construction of $G$, $C$ is a clique and $I$ is an independent set, and hence $V(G)=C\cup I$ is a clique-independent set partition of $V(G)$. Thus, $G(C\cup I,E)$ is a split graph. Further, by the construction of $G$, it is clear that for $y_{w}\in I$, $N_{G}(y_{w})=\{z_{(x_{1},1)},z_{(x_{2},2)},\ldots,z_{(x_{r},r)}\}$ if $w=(x_{1},x_{2},\ldots,x_{r})$. Hence, $deg_{G}(y_{w})=r$ for every $y_{w}\in I$. Thus, $G$ is an $I_{r}$-split graph. Next, we prove the guarantee of Construction 4. The following claim is used in the proof of guarantee of Construction 4.
@@
On the contrary, assume that there exists an open packing $S$ in $G$ such that $S\cap C$ is non-empty and $|S|>1$. Let $u\in S\cap C$, and let $z\in S\setminus\{u\}$. Since $r\geq 3$ and $q\geq 1$, $|C|\geq 3$. This implies that $C$ is a clique of size at least three in $G$ and so any two vertices in $C$ have a common neighbour in $G$. This says that $z\notin C$. So, $z\in I$, i.e., $z=y_{w}$ for some $w\in M$. Since $r\geq 3$ and $deg_{G}(y_{w})=r$ with $N_{G}(y_{w})\subseteq C$, $(z=)y_{w}\in I$ is adjacent to at least two vertices other that $u$ in $C$. This implies that $N_{G}(y_{w})\cap N_{G}(u)\neq\emptyset$, a contradiction to $u,z(=y_{w})\in S$ and $S$ is an open packing in $G$.
@@
Given an $I_{r}$-split graph $G$ and a vertex subset $S$ of $G$, it can be tested in polynomial time whether $N_{G}(u)\cap N_{G}(v)$ is empty for every distinct $u,v\in S$. Hence, Open Packing is in the class NP on $I_{r}$-split graphs. Also, Construction 4 and the fact that $r$-Dimensional Matching is NP-complete for $r\geq 3$ [12] implies that Open Packing is NP-complete on $I_{r}$-split graphs for $r\geq 3$. ∎
The following construction is useful to show that Open Packing is polynomial time solvable on a superclass of $I_{1}$-split graphs and $I_{2}$-split graphs. Thus completing our dichotomy result on $I_{r}$-split graphs.
@@
Guarantee: For $k\geq 3$, $G$ has an open packing of size $k$ if and only if $G^{\prime}$ has a matching of size $k$.
@@
Let $J=\{v\in I\,:\,deg_{G}(v)=1\}$. Then, the vertex set and the edge set of $G^{\prime}$ are $V^{\prime}=C\cup J$ and $E^{\prime}=\{e_{v}\,:\,v\in I\}$. Note that $|V^{\prime}|\leq n$ and $|E^{\prime}|\leq n$. Hence, the graph $G^{\prime}$ can be constructed in $O(n)$ time. Also, note that the graph $G^{\prime}$ may have parallel edges. Next, we prove the guarantee of the above construction.
@@
Proof of Guarantee.
We prove the guarantee of Construction 5 in two cases.
Case 1: $G$ has an open packing $S$ such that $|S|=k\geq 3$.
Since $S$ is an open packing of size at least three in $G$, $S\subseteq I$ by Claim 9. Let $F_{S}=\{e_{v}\in E^{\prime}\,:\,v\in S\}$. We show that $F_{S}$ is a matching in $G^{\prime}$. On the contrary, assume that $F_{S}$ is not a matching in $G^{\prime}$. Then, there exist two distinct edges $e_{v},e_{y}\in F_{S}$ such that $V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})\neq\emptyset$. Let $u\in V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})$, i.e., $e_{v},e_{y}\in E_{G^{\prime}}(u)$. Then, by the construction of $G^{\prime}$, $u\in C\cup J$ and $v,y\in I$. Firstly, we show that $u\notin J$. If not, $u\in J$. Then, $deg_{G}(u)=1$ and so by Step 3 of Construction 5, $E_{G^{\prime}}(u)=\{e_{u}\}$. This implies that $e_{v}=e_{u}=e_{y}$, a contradiction to $v\neq y$. So, $u\notin J$. Hence, $u\in C$, and so by the construction of $G^{\prime}$, $uv,uy\in E(G)$. This implies that $u\in N_{G}(v)\cap N_{G}(y)$, a contradiction to $v,y\in S$ and $S$ is an open packing in $G$. So, $F_{S}$ is a matching of size $k$ in $G^{\prime}$.
@@
Let $S_{F}=\{v\in I\,:\,y_{v}\in F\}$. In this case, we prove that $S_{F}$ is an open packing in $G$. On the contrary, assume that $S_{F}$ is not an open packing in $G$. Then, there exist vertices $v,y\in S_{F}\subseteq I$ such that $N_{G}(v)\cap N_{G}(y)\neq\emptyset$. Let $u\in(N_{G}(v)\cap N_{G}(y))\subseteq C$. This implies that $u\in V(G^{\prime})=C\cup J$. Also, by the construction of $G^{\prime}$, $u\in V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})$, a contradiction to $e_{v},e_{y}\in F$ and $F$ is a matching in $G^{\prime}$. Hence, $S_{F}$ is an open packing in $G$. ∎
Note that the above algorithm solves the case when $\rho^{o}(G)\geq 3$ and we use Lemma 4 to complete the case $\rho^{o}(G)\leq 2$.
@@
A maximum open packing in a split graph $G(C\cup I,E)$ with $1\leq deg_{G}(v)\leq 2$ for every $v\in I$ can be found in $O(n^{3})$ time.
@@
Let $G(C\cup I,E)$ be a split graph with $deg_{G}(v)\in\{1,2\}$ for every $v\in I$. Then, compute a maximum matching of $G^{\prime}$ constructed in Construction 5. If $G^{\prime}$ has a maximum matching $F$ of size at least three, then by the guarantee of Construction 5, $\rho^{o}(G)=|F|$. Also, by Case 2 in the proof of guarantee of Construction 5, $S_{F}=\{v\in I\,:\,y_{v}\in F\}$ is a maximum open packing in $G$. Note that the graph $G^{\prime}$ can be constructed in $O(n+m)$ time, where $m$ denotes the number of edges in $G$. Hence, the maximum matching $F$ of $G^{\prime}$ can be computed in $O(m\sqrt{n})$ time [15] because $|V(G^{\prime})|$ and $|E(G^{\prime})|$ are of $O(n)$. Thus, a maximum open packing $S_{F}$ of $G$ when $\rho^{o}(G)\geq 3$ can be computed in $O(m\sqrt{n})$ time. Else if, the size of a maximum matching in $G^{\prime}$ is at most two, then again by the guarantee of Construction 5, $\rho^{o}(G)\leq 2$. So, in such a case, to find an optimal open packing, it is enough to examine all the $\binom{n}{2}$ two-sized subsets of $G$ to know whether an open packing of size two exists. By Lemma 4, it takes at most $O(n^{3})$ time to test whether an open packing of size two exists. If one exists, then $\rho^{o}(G)=2$. Otherwise, no open packing of size two exist, and so $\rho^{o}(G)=1$ and every vertex in $G$ constitutes a maximum open packing of $G$. ∎
@@ Appendix A Open Packing
We start with giving some simple clarifications to get a clear understanding of our proof. We work based on the adjacency list of the graph $G$. We say that a vertex $u\in V(G)$ is visited by the algorithm if the vertex is chosen by at least one of the while loops during some stage of the algorithm. We call the while loop that is performed based on whether $U$ is empty or not as the external while loop. We refer to the iterations performed by the external while loop as the iterations of the algorithm. The while loop that is performed based on the set $W$ is called the internal while loop.
We begin with the proof for the termination of the algorithm. Note that if the algorithm visits some vertex that already marked, then the algorithm will terminate. Else, the algorithm visits no marked vertex in any iteration of the algorithm. But, since $W$ is a finite set for every iteration of the external while loop, the internal while loop iterates only a finite number of times, and hence every iteration of the external while loop will terminate. Also, since the size of $U$ is finite, the number of iterations performed by the algorithm is finite, and hence the algorithm would terminate once $U$ becomes empty.
Next, we prove the correctness of Algorithm 1. Let $S=\{u_{1},u_{2},\ldots,u_{k}\}$ and W.L.O.G., assume that the algorithm chooses the vertices in $S$ in the increasing order of its suffices. We prove the correctness of the algorithm in two cases based on whether the input vertex subset $S$ is an open packing in $G$ or not.
@@
Since every vertex in $V(G)$ is unmarked during the initialization process, no vertex in $N_{G}(u_{1})$ might be marked while examining them. For $2\leq i\leq k$, the vertices of $G$ marked in any of the previous iterations of the algorithm are in $\cup_{s=1}^{i-1}N_{G}(u_{s})$. Since $S$ is an open packing, $N_{G}(u_{i})\cap N_{G}(u_{s})=\emptyset$ for every $s<i$. These statements imply that no vertex in $N_{G}(u_{i})$ is marked before the $i^{th}$ iteration. So, the algorithm will never encounter a marked vertex, and hence the algorithm will return that ‘$S$ is an open packing in $G$’. Thus, the algorithm correctly determines the yes instances of the open packing sets of a graph.
@@
Since $S$ is not an open packing in $G$, there exist $i,s$ such that $1\leq i<s\leq k$ and $N_{G}(u_{i})\cap N_{G}(u_{s})\neq\emptyset$. W.L.O.G., assume that $u_{s}$ is the least indexed vertex such that there exists a vertex $u_{i}$ with $i<s$ and $N_{G}(u_{i})\cap N_{G}(u_{s})\neq\emptyset$. Note that $S^{\prime}=\{u_{1},\ldots,u_{s-1}\}$ is an open packing in $G$ by the choice of $s$. Hence, by the arguments in Case 1, the algorithm does not end before examining $N_{G}(u_{s})$. Let $N_{G}(u_{i})\cap N_{G}(u_{s})=\{w_{1},w_{2},\ldots,w_{p}\}$ for some $p\leq n$. Note that every vertex in $N_{G}(u_{i})$ is marked during the $i^{th}$ iteration of the external while loop. So, the algorithm will encounter a marked vertex ($w_{t}$ for some $1\leq t\leq p$) while exploring $N_{G}(u_{s})$ during the $s^{th}$ iteration. Hence, the algorithm will return ‘$S$ is not an open packing in $G$’. Thus, the algorithm correctly determines the no instances of the open packing sets of a graph.
@@
Next, we prove that the algorithm runs in $O(n)$ time. To prove this we show that the size of the list of vertices visited by the algorithm is at most $12n+$ 1. Let $Q=(v_{1},v_{2},\ldots,v_{r})$ for some $r\in\mathbb{N}$. Then, it is clear that $Q$ is the list of vertices visited by the internal while loop (i.e., every vertex $v_{j}$ (for $1\leq j\leq r$) is visited as a neighbour of $u_{i}$ for some $i\in\{1,2,\ldots,k\}$) before the termination of the algorithm. We claim that $v_{j}\neq v_{\ell}$ for every $j,\ell\in\{1,2,\ldots,r-1\}$ and $j\neq\ell$. On the contrary, assume that there exist $j,\ell\in\{1,2,\ldots,r-1\}$ such that $j\neq\ell$ but $v_{j}=v_{\ell}=u$ for some $u\in V(G)$. W.L.O.G., assume that $\ell$ is the least integer in $\{1,2,\ldots,r-1\}$ such that there exists $j<\ell$ with $v_{j}=v_{\ell}$. Then, the vertex $u(=v_{j})$ is marked by the algorithm when the internal while loop visits the $j^{th}$ vertex in $Q$. Also, by the assumption on $\ell$, every vertex in $(v_{1},\ldots,v_{\ell-1})$ is distinct, and hence the algorithm does not terminate without visiting $\ell^{th}$ vertex of $Q$. But, when the $\ell^{th}$ vertex $u=v_{j}$ of the array $Q$ is visited by the internal while loop, the algorithm would terminate as the vertex is already marked while visiting the $j^{th}$ vertex ($u$) in $Q$, a contradiction to the assumption that $r>\ell$ and the internal while loop visits $r$ vertices in $Q$ before the algorithm terminates. So, $v_{j}\neq v_{\ell}$ for every $j,\ell\in\{1,2,\ldots,r-1\}$ and $j\neq\ell$. This implies that the cardinality of the set $V^{\prime}=\{v_{1},v_{2},\ldots,v_{r-1}\}$ is $r-1$. Since $V^{\prime}\subseteq V(G)$, $r-1=|V^{\prime}|\leq n$. Hence, $r\leq n+1$, i.e., the size of the list of vertices visited by the internal while loop is at most $n+1$. Note that for $Q=(v_{1},v_{2},\ldots,v_{r})$ and $S=\{u_{1},u_{2},\ldots,u_{k}\}$, the list of vertices visited by the algorithm is a permutation of some subarray of $(v_{1},v_{2},\ldots,v_{r},u_{1},u_{2},\ldots,u_{k})$. We say subarray because the algorithm may terminate before visiting some vertices in $S$. Hence, the size of the list of vertices visited by the algorithm is at most $|S|+|Q|\leq 2n+1$. Also, note that the edges of $G$ are visited only during the internal while loop (for neighbourhood testing). If $E^{\prime}=(e_{1},e_{2},\ldots,e_{q})$ is the list of edges visited by the algorithm with the suffixes denoting the order in which the edges are visited, then by the construction of $Q$, it is clear that for every $e_{i}$ ($1\leq i\leq q$) visited by the algorithm exactly one vertex ($v_{i}$) in $Q$ visited. Since the size of $Q$ is at most $n+1$, $q\leq n+1$. Hence, the algorithm runs in $3n+2=O(n)$ time. ∎

### v2

@@ 4.2 $I_{r}$-split graphs
Figure 9: (a) An instance $(X_{1},X_{2},X_{3},X_{4},M)$ of $4$-Dimensional Matching and (b) An $I_{r}$-split graph $G(C\cup I,E)$ corresponding to the instance $(X_{1},X_{2},X_{3},X_{4},M)$ produced using Construction 4. The edges between the vertices of $C$ aren’t given in the figure for clarity. Note that for every $X_{i}=\{z_{i},u_{i},v_{i}\}$ in the instance of $4$-Dimensional Matching, three vertices $z_{(x_{i},i)},z_{(u_{i},i)}$ and $z_{(v_{i},i)}$ are created in the graph $G$.
An example of Construction 4 is given in Fig. 9. The vertex and edge sets of the graph $G$ are $V(G)=C\cup I$, where $I=\{y_{w}\,:\,w\in M\}$ and $E(G)=\{xx^{\prime}\,:\,x,x^{\prime}\in C\}\cup\{z_{(x,i)}y_{w}\,:\,x\in X_{i},w\in M,\text{ and the }i^{th}\text{ coordinate of }w\text{ is }x\}$. Then, $|V(G)|=(r\cdot q)+|M|$ and $|E(G)|=r\cdot|M|+\binom{r\cdot q}{2}$. Hence, the graph $G$ can be constructed in polynomial time with respect to $r,q,$ and $|M|$. Note that by the construction of $G$, $C$ is a clique and $I$ is an independent set, and hence $V(G)=C\cup I$ is a clique-independent set partition of $V(G)$. Thus, $G(C\cup I,E)$ is a split graph. Further, by the construction of $G$, it is clear that for $y_{w}\in I$, $N_{G}(y_{w})=\{z_{(x_{1},1)},z_{(x_{2},2)},\ldots,z_{(x_{r},r)}\}$ if $w=(x_{1},x_{2},\ldots,x_{r})$. Hence, $deg_{G}(y_{w})=r$ for every $y_{w}\in I$. Thus, $G$ is an $I_{r}$-split graph. Next, we prove the guarantee of Construction 4. The following claim is used in the proof of guarantee of Construction 4.
@@
On the contrary, assume that there exists an open packing $S$ in $G$ such that $S\cap C$ is non-empty and $|S|>1$. Let $u\in S\cap C$, and let $z\in S\setminus\{u\}$. Since $r\geq 3$ and $q\geq 1$, $|C|\geq 3$. This implies that $C$ is a clique of size at least three in $G$, and so any two vertices in $C$ have a common neighbour in $G$. This says that $z\notin C$. So, $z\in I$, i.e., $z=y_{w}$ for some $w\in M$. Since $r\geq 3$ and $deg_{G}(y_{w})=r$ with $N_{G}(y_{w})\subseteq C$, $(z=)y_{w}\in I$ is adjacent to at least two vertices other than $u$ in $C$. This implies that $N_{G}(y_{w})\cap N_{G}(u)\neq\emptyset$, a contradiction to $u,z(=y_{w})\in S$ and $S$ is an open packing in $G$.
@@
Given an $I_{r}$-split graph $G$ and a vertex subset $S$ of $G$, it can be tested in linear time whether $S$ is an open packing in $G$ or not using Lemma 4. Hence, Open Packing is in the class NP on $I_{r}$-split graphs. Further, Construction 4 and the fact that $r$-Dimensional Matching is NP-complete for $r\geq 3$ [12] implies that Open Packing is NP-complete on $I_{r}$-split graphs for $r\geq 3$. ∎
The following construction is used to show that Open Packing is polynomial time solvable on a superclass of $I_{1}$-split graphs and $I_{2}$-split graphs. Thus, completing our dichotomy result on $I_{r}$-split graphs.
@@
Guarantee: For $k\in\mathbb{N}$, $G$ has an open packing $S$ such that $S\subseteq I$ and $|S|=k$ if and only if $G^{\prime}$ has a matching of size $k$.
@@
Let $J=\{v\in I\,:\,deg_{G}(v)=1\}$. Then, the vertex set and the edge set of $G^{\prime}$ are $V^{\prime}=C\cup J$ and $E^{\prime}=\{e_{v}\,:\,v\in I\}$. Note that $|V^{\prime}|\leq n$ and $|E^{\prime}|\leq n$. Since $G^{\prime}$ is constructed based on $N_{G}(v)$ for every $v\in I$, it can be constructed in $O(n+m)$ time. Also, note that the graph $G^{\prime}$ may have parallel edges. Next, we prove the guarantee of the above construction.
@@
Proof of Guarantee of Construction 5.
We prove the guarantee of the construction in two cases.
Case 1: $G$ has an open packing $S$ such that $|S|=k$ and $S\subseteq I$.
Let $F_{S}=\{e_{v}\in E^{\prime}\,:\,v\in S\}$. We show that $F_{S}$ is a matching in $G^{\prime}$. On the contrary, assume that $F_{S}$ is not a matching in $G^{\prime}$. Then, there exist two distinct edges $e_{v},e_{y}\in F_{S}$ such that $V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})\neq\emptyset$. Let $u\in V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})$, i.e., $e_{v},e_{y}\in E_{G^{\prime}}(u)$. Then, by the construction of $G^{\prime}$, $u\in C\cup J$ and $v,y\in I$. Firstly, we show that $u\notin J$. If not, $u\in J$. Then, $deg_{G}(u)=1$, and so, by Step 3 of Construction 5, $E_{G^{\prime}}(u)=\{e_{u}\}$. This implies that $e_{v}=e_{u}=e_{y}$, a contradiction to $v\neq y$. So, $u\notin J$. Hence, $u\in C$, and so by the construction of $G^{\prime}$, $uv,uy\in E(G)$. This implies that $u\in N_{G}(v)\cap N_{G}(y)$, a contradiction to $v,y\in S$ and $S$ is an open packing in $G$. So, $F_{S}$ is a matching of size $k$ in $G^{\prime}$.
@@
Let $S_{F}=\{v\in I\,:\,y_{v}\in F\}$. In this case, we prove that $S_{F}$ is an open packing in $G$. On the contrary, assume that $S_{F}$ is not an open packing in $G$. Then, there exist vertices $v,y\in S_{F}\subseteq I$ such that $N_{G}(v)\cap N_{G}(y)\neq\emptyset$. Let $u\in(N_{G}(v)\cap N_{G}(y))\subseteq C$. This implies that $u\in V(G^{\prime})=C\cup J$. Also, by the construction of $G^{\prime}$, $u\in V_{G^{\prime}}(e_{v})\cap V_{G^{\prime}}(e_{y})$, a contradiction to $e_{v},e_{y}\in F$ and $F$ is a matching in $G^{\prime}$. Hence, $S_{F}$ is an open packing in $G$ such that $S\subseteq I$ and $|S|=k$. ∎
The following lemma and Construction 5 help us compute a maximum open packing in $I_{1}$-split graphs and $I_{2}$-split graphs.
Lemma 13 ([15]).
A maximum matching in a graph with $n$ vertices and $m$ edges can be found in $O(m\sqrt{n})$ time.
@@
A maximum open packing in a split graph $G(C\cup I,E)$ with $1\leq deg_{G}(v)\leq 2$ for every $v\in I$ can be found in $O(\max\{n+m,n^{1.5}\})$ time.
@@
Let $G(C\cup I,E)$ be a split graph with $deg_{G}(v)\in\{1,2\}$ for every $v\in I$. Also, let $F$ be a maximum matching of the graph $G^{\prime}$ constructed in Construction 5, i.e., $\alpha^{\prime}(G^{\prime})=|F|$. Then, we find the open packing number of $G$ using $F$ in four parts.
Part 1: If $\alpha^{\prime}(G^{\prime})\geq 3$, then $S_{F}=\{v\in I\,:\,e_{v}\in F\}$ is a maximum open packing in $G$.
Note that $S_{F}$ is an open packing in $G$ by Case 2 in the proof of Guarantee of Construction 5. So, $\rho^{o}(G)\geq|S_{F}|=\alpha^{\prime}(G)\geq 3$. Hence, by Claim 9, every maximum open packing in $G$ is a subset of $I$. Let $S$ be a maximum open packing in $G$. Then, by Case 1 in the proof of Guarantee of Construction 5, $F_{S}=\{e_{v}\in E^{\prime}\,:\,v\in S\}$ is a matching in $G^{\prime}$. So, $\alpha^{\prime}(G^{\prime})\geq|F_{S}|=|S|=\rho^{o}(G)\geq|S_{F}|=\alpha^{\prime}(G^{\prime})$. Hence, $\rho^{o}(G)=\alpha^{\prime}(G^{\prime})$, and $S_{F}$ is a maximum matching in $G^{\prime}$.
Part 2: If $\alpha^{\prime}(G^{\prime})=2$, then $S_{F}=\{v\in I\,:\,e_{v}\in F\}$ is a maximum open packing in $G$.
If $\rho^{o}(G)\geq 3$, then by Claim 9 and Construction 5, $\alpha^{\prime}(G^{\prime})\geq 3$, a contradiction to the assumption that $\alpha^{\prime}(G^{\prime})=2$. So, $\rho^{o}(G)\leq 2$. By Case 2 in the proof of Guarantee of Construction 5, $S_{F}$ is an open packing of size two in $G$. Hence, $\rho^{o}(G)=2$.
Part 3: If $\alpha^{\prime}(G^{\prime})=1$ and there exists a vertex of degree one in $G$, then $\rho^{o}(G)=2$.
Since $\alpha^{\prime}(G^{\prime})=1$, it is evident, by Claim 9 and Construction 5, that $\rho^{o}(G)\leq 2$. Let $v\in V(G)$ such that $deg_{G}(v)=1$. Then, $N_{G}[v]$ is an open packing of size two in $G$. Hence, $\rho^{o}(G)=2$.
Part 4: If $\alpha^{\prime}(G^{\prime})=1$ and every vertex in $G$ is of degree at least two, then $\rho^{o}(G)=1$.
On the contrary, assume that $G$ has an open packing $S$ of cardinality at least two. Since $\alpha^{\prime}(G^{\prime})=1$, there does not exist an open packing $S$ in $G$ such that $|S|\geq 2$ and $S\subseteq I$ by Construction 5. So, $S\cap C\neq\emptyset$. Let $u\in S\cap C$. Then, by Claim 9, $S\subseteq N_{G}[u]$. Let $v\in S\setminus\{u\}$. Then, by the assumption on Part 4 that every vertex in $G$ is of degree at least two, for any choice of $v$ in $V(G)\setminus\{u\}=(C\cup I)\setminus\{u\}$, $u$ and $v$ must have a common neighbour in $G$, a contradiction to $S$ being an open packing in $G$ with $u,v\in S$. So, $\rho^{o}(G)=1$.
Parts 1-4 complete the arguments for finding a maximum open packing in $G$. Next, we look at the complexity of computing it. Note that
  • (a)
    the graph $G^{\prime}$ can be constructed in $O(n+m)$ time,
  • (b)
    a maximum matching in $G^{\prime}$ can be found in $O(|E(G^{\prime})|\sqrt{V(G^{\prime})})=O(n^{1.5})$ time using Lemma 13, and
  • (c)
    degree of every vertex in $G$ can be found in $O(n+m)$ time.
By (a)-(c) and Parts 1-4, a maximum open packing in $G$ can be found in $O(\max\{n+m,n^{1.5}\})$ time. ∎
@@ Appendix A Open Packing
We start by giving some simple clarifications to get a clear understanding of our proof. We use the adjacency list of the graph $G$ as the data structure. We say that a vertex $u\in V(G)$ is visited by the algorithm if the vertex is chosen by at least one of the while loops during some stage of the algorithm. We call the while loop that is performed based on whether $U$ is empty or not the external while loop. We refer to the iterations performed by the external while loop as the iterations of the algorithm. The while loops that are performed based on the set $W$ are called the internal while loops.
We begin with the proof for the termination of the algorithm. Note that if the algorithm visits some vertex that is already marked, then the algorithm will terminate. Otherwise, the algorithm visits no marked vertex in any iteration of the algorithm. However, since $W$ is a finite set for every iteration of the external while loop, the internal while loop iterates only a finite number of times. Hence, every iteration of the external while loop will terminate. Also, since the size of $U$ is finite, the number of iterations performed by the algorithm is finite, and hence, the algorithm would terminate once $U$ becomes empty.
Next, we prove the correctness of Algorithm 1. Let $S=\{u_{1},u_{2},\ldots,u_{k}\}$. W.L.O.G., assume that the algorithm chooses the vertices in $S$ in the increasing order of their suffices. We prove the correctness of the algorithm in two cases based on whether the input vertex subset $S$ is an open packing in $G$ or not.
@@
Since every vertex in $V(G)$ is unmarked during the initialization process, no vertex in $N_{G}(u_{1})$ might be marked while examining them. For $2\leq i\leq k$, the vertices of $G$ that are marked during the previous iterations of the algorithm are in $\bigcup\limits_{s=1}^{i-1}N_{G}(u_{s})$. Since $S$ is an open packing, $N_{G}(u_{i})\cap N_{G}(u_{s})=\emptyset$ for every $s<i$. These statements imply that no vertex in $N_{G}(u_{i})$ is marked before the $i^{th}$ iteration. So, the algorithm will never encounter a marked vertex, and hence the algorithm will return ‘$S$ is an open packing in $G$’. Thus, the algorithm correctly determines the yes instances of the open packing sets of a graph.
@@
Since $S$ is not an open packing in $G$, there exist $i,s$ such that $1\leq i<s\leq k$ and $N_{G}(u_{i})\cap N_{G}(u_{s})\neq\emptyset$. W.L.O.G., assume that $u_{s}$ is the least indexed vertex such that there exists a vertex $u_{i}$ with $i<s$ and $N_{G}(u_{i})\cap N_{G}(u_{s})\neq\emptyset$. Note that $S^{\prime}=\{u_{1},\ldots,u_{s-1}\}$ is an open packing in $G$ by the choice of $s$. Hence, by the arguments in Case 1, the algorithm does not end before examining $N_{G}(u_{s})$. Let $N_{G}(u_{i})\cap N_{G}(u_{s})=\{w_{1},w_{2},\ldots,w_{p}\}$ for some $p\leq n$. Note that every vertex in $N_{G}(u_{i})$ is marked during $i^{th}$ iteration of the external while loop. So, the algorithm will encounter a marked vertex ($w_{t}$ for some $1\leq t\leq p$) while exploring $N_{G}(u_{s})$ during the $s^{th}$ iteration. Hence, the algorithm will return ‘$S$ is not an open packing in $G$’. Thus, the algorithm correctly determines the no instances of the open packing sets of a graph.
@@
Next, we prove that the algorithm runs in $O(n)$ time. To prove this, we show that the size of the list of vertices visited by the algorithm is at most $12n+$ 1. Let $Q=(v_{1},v_{2},\ldots,v_{r})$ for some $r\in\mathbb{N}$. Then, it is clear that $Q$ is the list of vertices visited by the internal while loop (i.e., every vertex $v_{j}$ (for $1\leq j\leq r$) is visited as a neighbor of $u_{i}$ for some $i\in\{1,2,\ldots,k\}$) before the termination of the algorithm. We claim that $v_{j}\neq v_{\ell}$ for every $j,\ell\in\{1,2,\ldots,r-1\}$ and $j\neq\ell$. On the contrary, assume that there exist $j,\ell\in\{1,2,\ldots,r-1\}$ such that $j\neq\ell$ but $v_{j}=v_{\ell}=u$ for some $u\in V(G)$. W.L.O.G., assume that $\ell$ is the least integer in $\{1,2,\ldots,r-1\}$ such that there exists $j<\ell$ with $v_{j}=v_{\ell}$. Then, the vertex $u(=v_{j})$ is marked by the algorithm when the internal while loop visits the $j^{th}$ vertex in $Q$. Also, by the assumption on $\ell$, every vertex in $(v_{1},\ldots,v_{\ell-1})$ is distinct, and hence the algorithm does not terminate without visiting $\ell^{th}$ vertex of $Q$. But, when the $\ell^{th}$ vertex $u=v_{\ell}$ of the array $Q$ is visited by the internal while loop, the algorithm would terminate as the vertex is already marked while visiting the $j^{th}$ vertex ($u$) in $Q$, a contradiction to the assumption that $r>\ell$ and the internal while loop visits $r$ vertices in $Q$ before the algorithm terminates. So, $v_{j}\neq v_{\ell}$ for every $j,\ell\in\{1,2,\ldots,r-1\}$ and $j\neq\ell$. This implies that the cardinality of the set $V^{\prime}=\{v_{1},v_{2},\ldots,v_{r-1}\}$ is $r-1$. Since $V^{\prime}\subseteq V(G)$, $r-1=|V^{\prime}|\leq n$. Hence, $r\leq n+1$, i.e., the size of the list of vertices visited by the internal while loop is at most $n+1$. Note that for $Q=(v_{1},v_{2},\ldots,v_{r})$ and $S=\{u_{1},u_{2},\ldots,u_{k}\}$, the list of vertices visited by the algorithm is a permutation of some subarray of $(v_{1},v_{2},\ldots,v_{r},u_{1},u_{2},\ldots,u_{k})$. We say subarray because the algorithm may terminate before visiting some vertices in $S$. Hence, the size of the list of vertices visited by the algorithm is at most $|S|+|Q|\leq 2n+1$. Also, note that the edges of $G$ are visited only during the internal while loop (for neighborhood testing). If $E^{\prime}=(e_{1},e_{2},\ldots,e_{q})$ is the list of edges visited by the algorithm with the suffices denoting the order in which the edges are visited, then by the construction of $Q$, it is clear that for every $e_{i}$ ($1\leq i\leq q$) visited by the algorithm, exactly one vertex ($v_{i}$) in $Q$ was visited. Since the size of $Q$ is at most $n+1$, $q\leq n+1$. Hence, the algorithm runs in $3n+2=O(n)$ time. ∎